Find 3 consecutive even numbers where the product of the smaller two numbers is 40 less than the square of the largest number.
Answer by mananth(16946)

Find 3 consecutive even numbers where the product of the smaller two numbers is 40 less than the square of the largest number.
..
x,x+2,x+4 be the even numbers
x(x+2)=(x+4)^2-40
x^2+2x=x^2+8x+16-40
reorganise the terms and constants
x^2 cancels off
8x-2x=40-16
6x=24
/6
x=4
the numbers are 4,6,8
m.ananth@hotmail.ca